A .4 kilogram ball is rolled off a 1 meter high level table. If the ball lands in line with a spring of 50 N/M and the spring compresses exactly .5 meters, how much time did our ball spend in the air..what horizontal displacement did it undergo.. what was the kinetic energy of the ball before it reached the spring.. and what was the initial velocity of the spring?
Answer #1

time spent in the air is found by the displacement formula:

h = 1/2 g t^2

=> 1 = 1/2 g t^2

so t = sqrt (2/g) = 0.452 s

________________________________

now find v of the ball when it just lefts the table :

=> (KE + PE)1 = 1/2 k x^2

=> 1/2 (0.4) v^2 + 0.4(9.8)(1.5) = 25 (0.5)^2

=> v = 1.36 m/s

so the horizontal distance is :

dx = v * t = 1.36 * 0.452 = 0.615 m

________________________________________

the kinetic energy of the ball before it reached the spring is :

KE = 1/2 k x^2 = 25 (0.25) = 6.25 J

_______________________________________

the initial velocity of the spring :

=> KE = 1/2 m v^2

=> 6.25 = 0.2 v^2

=> v = 5.59 m/s